SEMI-EXPERIMENTAL (r_{s}/r_{e}) STRUCTURES FOR THE HEAVY ATOM BACKBONES OF TWO MOD-ERATELY LARGE MOLECULES OBTAINED FROM MICROWAVE SPECTROSCOPY AND QUANTUM CHEMICAL CALCULATIONS

\underline{NORMAN C. CRAIG}, Department of Chemistry and Biochemistry, Oberlin College, Oberlin, OH 44074; ALBERTO LESARRI, Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47011 Valladolid, Spain; EMILIO J. COCINERO, Departamento de Química Física, Facultad de Ciencia y Tecnología, Universidad del País Vasco, Ap. 644, E-48080 Bilbao, Spain; JENS-UWE GRABOW, Institut für Physikalische Chemie und Elektrochemie, Gottfried-Wilhelm-Leibniz-Universität Hannover, Callinstrasse 3A, D30167 Hannover, Germany.

From recent microwave investigations of 1-methyl-4-piperidone^{a} and tropinone^{b} ground state rotational constants are available for the equatorial conformers of the normal species and the isotopologues with single substitution of all the heavy atoms. Vibration-rotation constants (alphas) were computed with Gaussian 03 (G03) for the B3LYP/cc-pVTZ model and used to convert ground state rotational constants into equilibrium rotational constants. Using the Kraitchman equations (r_{s} method), the equilibrium (r_{e}) Cartesian coordinates were determined for all the heavy atoms in the principal axis framework. Equilibrium bond lengths and bond angles are compared with those computed with the B3LYP/cc-pVTZ model. We have compared the ground state rotational constants computed with G03, after scaling by factors based on the normal species, with observed values. The agreement is within 0.1% for the full set of constants (0.04% for methyl-piperidone and 0.1% for tropinone). This agreement between experiment and theory is so good that it seems possible to use calculated ground state rotational constants in place of experimental ones for determining r_{s}/r_{e} structures for organic molecules of this size.

aL. Evangelisti, A. Lesarri, M. Jahn, E. Cocinero, W. Caminati, J.-U. Grabow J. Phys. Chem. A, submitted.

bE. J. Concinero, A. Lesarri, P. Ecija, J.-U. Grabow, J. A. Fernandez, F. Castano PCCP 12, 6076-6083 (2010).